What is Ohm's law in vector form.........??????????????

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Answer :

Simply ohm’s law is:
V = IR
where V is the potential difference, I is the current and R is the resistance.
Also R = pl/A
where pis the resistivity which is material dependent parameter.
Now current density J is a vector quantity given by Jz = I/A if the conductor is aligned along z axis.
Substituting R in V = IR,
V = I * pl/A
Putting value of I/A = J
V = plJ
Also V/l is E which is a vector quantity , so
E = pJ
Or J = σE

This is expressed in vector form.

Thanks.